# Universal consistency of the $k$-NN rule in metric spaces and Nagata dimension. III

### v1

@@ Abstract.
We prove the last remaining implication allowing to claim the equivalence of the following conditions for a complete separable metric space $X$:
@@
    The $k$-nearest neighbour classifier is (weakly) universally consistent in $X$,
@@
    $X$ is sigma-finite dimensional in the sense of Nagata.
@@
$k$-NN classifier, universal consistency, Nagata dimension, sigma-finite dimensional metric spaces, Lebesgue–Besicovitch property
@@ 1. Introduction
This is the third and last article in a series [4, 11], but it is to a large degree self-contained. The main new result is the following.
{thrm}
Namely, let $(k_{n})$ be an increasing sequence of natural numbers with $k_{n}\in\omega(1)\cap o(n)$. There exist a probability measure $\mu$ on $\Omega$ and a measurable regression function $\eta\colon\Omega\to\{0,1\}$, with $P[\eta(X)=0]=1/2=P[\eta(X)=1]$, as well as a subsequence $(n_{i})$, such that the predictions of the $k$-NN classifier corresponding to the values $n=n_{i}$ and $k=k_{n_{i}}$ converge in probability to the identical value $1$ as $i\to\infty$.
The result holds under any strategy of distance tie-breaking. The proof in essence extends the construction of the measure due to Preiss [17] from the Hilbert space $\ell^{2}$ to any complete separable metric space that is not sigma-finite dimensional in the sense of Nagata, as defined by Preiss elsewhere [18].
There are two consequences of the above result. First, we can now claim the validity of the conjecture put forward in [4].
@@
    The $k$-NN classifier is (weakly) universally consistent in $\Omega$.
@@
    $\frac{1}{\mu(B_{r}(x))}\int_{B_{r}(x)}f(x)\,d\mu(x)\to f(x).$
@@
(Here we assume, like in [3], the uniform random distance tie-breaking of the $k$-NN classifier in case of distance ties.)
@@
The second consequence is that the weak universal consistency of the $k$-NN rule indeed requires the strong, rather than weak, Lebesgue–Besicovitch differentiation property. Now we can indicate an example of a complete separable metric space and a probability measure on it which satisfies the weak Lebesgue–Besicovitch differentiation property, yet the $k$-NN classifier is not universally consistent. This is the Heisenberg group equipped with a Cygan–Koranyi metric, see e.g. [9, 19, 12] or a relevant discussion and references in [11]. The Heisenberg group is known to have finite metric dimension in the sense of de Groot [5], which, according to Assouad and Quentin de Gromard [1], implies it has the weak Lebesgue–Besicovitch differentiation property for every measure. In [11], we assumed the validity of the result from [3] that the weak Lebesgue–Besicovitch differentiation property suffices for the weak consistency of the $k$-NN classifier, and so we erroneously claimed that the $k$-NN classifier is universally consistent in the Heisenberg group. Our Theorem 1 implies this is not true, as the Heisenberg group is not sigma-finite dimensional in the sense of Nagata [9, 19].
We set things up in Section 2. In Section 3 we discuss the Nagata dimension. The main technical construction of a measure is given in Sect. 4, followed by the proof of Theorem 1 in Sect. 5. Finally, in Section 6 we observe that the weak Lebesgue–Besicovitch property does not in general suffice for the $k$-NN classifier to be consistent.
@@
This note was prompted by a private message [2] from the first author of the paper [3] correcting the statement of their main theorem. For this, I am most grateful to Frédéric Cérou, as well as to László Györfi, Aryeh Kontorovich, Harro Walk, and Roi Weiss for stimulating discussions, without which this paper would not have appeared.
@@ 2.1. General setting
We are studying the $k$ nearest neighbour rule for supervised binary classification, working in a complete separable metric space $\Omega$ equipped with a Borel probability measure $\mu$ and a measurable regression function $\eta\colon\Omega\to[0,1]$. (We prefer to denote the “domain” by $\Omega$, reserving the letter $X$ for its random element.) Elements of the random path are sampled in an i.i.d. way following the law $\mu$ and then every point $x$ is labelled 0 or $1$ according to the Bernoulli law with probability of success $\eta(x)$.
@@
We are interested in the conditons assuring the universal consistency of the $k$-NN classifier, that is, the property that for any pair $\mu,\eta$ as above, as $n,k\to\infty$ so that $k/n\to 0$, the expected misclassification error converges to the Bayes (the smallest possible) error as the size of the sample goes to infinity.
@@ 2.3. Implication (2) $\Rightarrow$ (1)
$\lim_{r\downarrow 0}\frac{1}{\mu(B_{r}(x))}\int_{B_{r}(x)}f(x)\,d\mu(x)=f(x).$ (1)
@@
$\frac{1}{\mu(B_{r}(x))}\int_{B_{r}(x)}\eta(x)\,d\mu(x)\to\eta(x)\mbox{ as }r\downarrow 0,$ (2)
@@
$\mu\left\{x\in\Omega\colon\left|\frac{1}{\mu(B_{r}(x))}\int_{B_{r}(x)}\eta(x)\,d\mu(x)-\eta(x)\right|>{\epsilon}\right\}\to 0\mbox{ when }r\downarrow 0.$
The original theorem in [3] stated that for every complete separable metric space with probability measure $\mu$ satisfying the weak Lebesgue–Besicovitch differentiation property the $k$-NN classifier is consistent with regard to any $\mu$-measurable regression function. However, one of the authors (Frédéric Cérou) has kindly pointed out to us [2] that the very last assertion in the proof of the main result of [3], at the bottom of p. 345, requires the strong Lebesgue–Besicovitch property.
@@
[Cérou and Guyader [3]] Let $\Omega$ be a complete separable metric space equipped with a probability measure, $\mu$. If $(\Omega,\mu)$ satisfies the strong Lebesgue–Besicovitch differentiation property, then for any $\mu$-measurable regression function $\eta\colon\Omega\to[0,1]$ the $k$-NN classifier is consistent.

### v2

@@ Abstract.
We establish the last missing link allowing to describe those complete separable metric spaces $X$ in which the $k$ nearest neighbour classifier is universally consistent, both in combinatorial terms of dimension theory and via a fundamental property of real analysis. The following are equivalent:
@@
    The $k$-nearest neighbour classifier is universally consistent in $X$,
@@
    $X$ is sigma-finite dimensional in the sense of Jun-Iti Nagata.
@@
$k$-NN classifier, $1$-NN classifier, Bayes error, universal consistency, Nagata dimension, sigma-finite dimensional metric spaces, Lebesgue–Besicovitch differentiation property
@@ 1. Introduction
The $k$ nearest neighbour classification rule is arguably one of the oldest learning algorithms, steeped in real-life practices, and at the same time it is among the learning rules most subjected to rigorous theoretical analysis. It was historically the first learning rule whose universal consistency (in the Euclidean space) was established, in the highly influential 1977 work of Charles J. Stone [30]. (Recall that a learning rule is universally consistent if its expected learning error converges, under any distribution of the labelled data, to the optimal, Bayes, error as the size of the random sample increases.) There is at least one monograph entirely dedicated to the $k$ nearest neighbour method [5]. At the same time, we are still far from fully understanding the algorithm and there are serious theoretical challenges in the area still to be addressed. The present paper aims to resolve one of the most natural such challenges.
Ample empirical evidence suggests that the efficiency of the $k$-NN algorithm depends on the choice of the domain metric, which may or may not be Euclidean. There is an extensive literature devoted to the choice of a metric based on the data, see the surveys [19, 4]. At the same time, the universal consistency of the variable metric $k$-NN algorithms, such as the Large Margin Nearest Neighbour (LMNN) classifier [34], is still an open question. So one of the most natural problems to begin with would be to determine for which metrics on a given dataset is the $k$-NN classifier universally consistent.
Here a pivotal investigation belongs to Cérou and Guyader [7]. First, they have ruled out non-separable metric spaces via a heuristic argument, which was later formalized by Hanneke et al. [15] to lead to the following alternative: in a non-separable metric space, either the $k$-NN rule is inconsistent (assuming the existence of the so-called large cardinals), or the $k$-NN consistence completely reduces to the same property in separable subspaces. Thus, only separable metric spaces deserve investigation. And by adding the completeness condition we do not lose in generality (as every metric space can be completed), yet the study becomes technically simpler. Thus, we can and will naturally restrict ourselves to complete separable metric spaces.
The second achievement of the above paper [7] was to establish the link between real analysis in metric spaces with measure and the consistency of the $k$-NN classifier. Recall that a model of statistical learning in a domain (in our case, metric space) $X$ states that the data is generated in an i.i.d. fashion following a probability measure $\mu$, and the labels of datapoints $x$ are independent Bernoulli variables with variable probability of success $\eta(x)$, where $\eta$ is the regression function on $X$. The main result of [7] is that if the regression function $\eta$ satisfies the strong Lebesgue–Besicovitch differentiation property with regard to the measure $\mu$, then the $k$-NN classifier is consistent.
In real analysis, the strong Lebesgue–Besicovitch differentiation property (defined a few lines below) is of great importance in various contexts. For the metric spaces, it was thoroughly investigated in particular by David Preiss [28] who in particular has characterized the metric spaces possessing this property under all probability measures in terms of a certain combinatorial parameter of metric geometry, discovered by Nagata [23] in his topological investigations: the metric dimension.
Now we can claim the validity of the following result, where statistical learning, real analysis, and dimension theory all come together.
@@
    The $k$-NN classifier is universally consistent in $\Omega$.
@@
    $\frac{1}{\mu(B_{r}(x))}\int_{B_{r}(x)}f(x^{\prime})\,d\mu(x^{\prime})\to f(x)~~\mbox{when $r\to 0$}.$
@@
Here $B_{r}(x)$ denotes the open ball of radius $r$ around $x$. As always, we say that the $k$-NN classifier is universally consistent if for any distribution of labelled data, the misclassification error of the classifier on the i.i.d. labelled data $n$-samples converges to the smallest possible (Bayes) misclassification error when $n,k\to\infty$ and $k/n\to 0$. Here we assume, like in [7], the uniform random distance tie-breaking of the $k$-NN classifier in case of distance ties.
The last property, that of $\Omega$ being sigma-finite dimensional, is a combinatorial property of metric spaces as defined by Preiss [28] based on the work of Nagata [23, 24] in topological dimension theory. We devote the entire Section 3 to treating this property in great detail.
@@
[…]
{thrm}
Namely, let $k_{n}\to\infty$ be a sequence of natural numbers with $k_{n}/n\to 0$. There exist a Borel probability measure $\mu$ on $\Omega$ and a measurable regression function $\eta\colon\Omega\to\{0,1\}$, with $P[\eta(X)=0]=1/2=P[\eta(X)=1]$, such that the misclassification error of the $k$-NN classifier using $k_{n}$ nearest neighbours in a random i.i.d. $n$-sample following the law $\mu$ with $\eta$ as conditional probability of label $1$ does not converge in probability to the Bayes error (zero in this case).
The result holds under any strategy of distance tie-breaking. The measure $\mu$ built in the proof in essence extends the construction of a measure due to Preiss [27] from the Hilbert space $\ell^{2}$ to any complete separable metric space that is not sigma-finite dimensional in the sense of Nagata, as defined by Preiss elsewhere [28]. We present the proof in Section 4.
In addition to Theorem 1, another consequence of Theorem 1 is the observation that the universal consistency of the $k$-NN rule indeed requires the strong, rather than weak, Lebesgue–Besicovitch differentiation property. The original theorem of [7] affirmed the consistency of the $k$-NN algorithm under the assumption that the regression function satisfies the weak version of the Lebesgue–Besicovitch property, that is, the convergence in eq. (2) in measure, rather than almost everywhere. However, it was noted by one of the authors [6] that the proof implicitly uses the strong version of the property.
@@
And indeed, now we can indicate examples of complete separable metric spaces and a probability measures on them which satisfy the weak Lebesgue–Besicovitch differentiation property, yet the $k$-NN classifier is not universally consistent. One of them is a recent example by Käenmäki, Rajala and Suomala [16]. Another is the Heisenberg group equipped with the Carnot–Cygan–Koranyi metric, see e.g. [18, 29, 22] or a relevant discussion and references in [21]. The Heisenberg group is known to have finite metric dimension in the sense of de Groot [11], which, according to Assouad and Quentin de Gromard [3], implies it has the weak Lebesgue–Besicovitch differentiation property for every measure.
In [21], we assumed the validity of the result from [7] that the weak Lebesgue–Besicovitch differentiation property suffices for the weak consistency of the $k$-NN classifier, and erroneously claimed that the $k$-NN classifier is universally consistent in the Heisenberg group. Our Theorem 1 implies this is not true, as the Heisenberg group is not sigma-finite dimensional in the sense of Nagata [18, 29]. This is the content of Section 5.
Another interesting consequence of the main result, given in Section 6, is that every complete separable metric space of uncountable cardinality admits a uniformly equivalent metric under which the $k$-NN classifier is not universally consistent – which looks rather counter-intuitive in the particular case of the real line.
The equivalent conditions in Theorem 1 can be also expressed in terms of the $1$-nearest neighbour classifier. We are stating the result separately so as not to overload the original theorem. In the classical theory of nearest neighbour classifier in the Euclidean space, it is known that the error of the $1$-NN classifier, asymptotically, is at most twice as bad as the Bayes error [9, 30]. This property is characteristic of the sigma-finite dimensional metric spaces as well.
{thrm}
Let $X$ be a complete metric space. The following two conditions are equivalent between themselves and to all the conditions in Theorem 1:
  1. (4)
    Under every distribution of the labelled data on $X\times\{0,1\}$, the misclassification error of the $1$-nearest neighbour rule, asymptotically as the size of an i.i.d. labelled sample goes to infinity, does not exceed twice the Bayes error.
  2. (4)
    Under every distribution of the labelled data on $X\times\{0,1\}$ whose Bayes error is zero (that is, the regression function is deterministic), the $1$-nearest neighbour rule is consistent when the size of an i.i.d. labelled sample goes to infinity.
This result is established in the concluding Section 7.
This is the third and last article in a series [8, 21], but it is largely self-contained. We set things up in Section 2. In Section 3 we discuss the Nagata dimension, and the remaining Sections contain proofs.
@@
This note was prompted by a private message [6] from the first author of the paper [7] correcting the statement of their main theorem. For this, I am most grateful to Frédéric Cérou, as well as to László Györfi, Aryeh Kontorovich, Harro Walk, and Roi Weiss for stimulating discussions, without which this paper would not have appeared. The author is equally grateful to the anonymous journal referee for the preliminary report that has prompted a very serious reworking of the first version of the article.
@@ 2.1. General setting
The standard model for supervised binary classification in statistical learning theory starts with the domain, which is a measurable space, that is, a set, $\Omega$, equipped with a sigma-algebra of subsets $\mathcal{A}$. (We use throughtout the rest of the paper the letter $\Omega$, reserving the letter $X$ for its random element.) The theory is only developed to any extent in the particular case where $\Omega$ is a standard Borel space. In other words, there is a metric $\rho$ on $\Omega$ making it a complete separable metric space, and such that the sigma-algebra $\mathcal{A}$ consists of the Borel subsets, that is, $\mathcal{A}$ is the smallest sigma-algebra containing all the open sets. It is a fundamental result of descriptive set theory that any two standard Borel spaces of the same cardinality are isomorphic between themselves, and any such space of uncountable cardinality is isomorphic, as a measurable space, to the unit interval (Th. 15.6 in [17]). In our case, $\Omega$ will be a complete separable metric space, thus, in particular, a standard Borel space.
A classifier in a domain $\Omega$ is a measurable mapping $T\colon\Omega\to\{0,1\}$. A finite sample in $\Omega$ is a sequence $(x_{1},x_{2},\ldots,x_{n})$, $x_{i}\in\Omega$, $n\in{\mathbb{N}}_{+}$. We will consider the case of binary labels taking values 0 or $1$. A labelled sample is a finite sequence of elements of the labelled domain, $\Omega\times\{0,1\}$, written $\sigma=(x_{1},y_{1},\ldots,x_{n},y_{n})$, $x_{i}\in\Omega$, $y_{i}\in\{0,1\}$. A learning rule for $\Omega$ is a countable sequence $g=(g_{n})_{n=1}^{\infty}$ of mappings, associating to each labelled sample a classifier, $g_{n}(\sigma)\colon\Omega\to\{0,1\}$. Each $g_{n}$ can be viewed as a mapping
$g_{n}\colon\Omega^{n}\times\{0,1\}^{n}\times\Omega\to\{0,1\},$
and we request those mappings to be measurable (with regard to the product sigma-algebra).
It is further assumed that the labelled pairs $(X,Y)$ are random elements of $\Omega\times\{0,1\}$ following a certain (unknown) probability measure $\tilde{\mu}$. The samples are i.i.d. sequences of such random elements following the same law.
It is usually convenient to reinterpret $\tilde{\mu}$ in a different way. Denote $\pi\colon\Omega\times\{0,1\}\to\Omega$ the first coordinate projection, $\pi(x,y)=x$. The direct image of $\tilde{\mu}$ under this projection is a probability measure on $\Omega$, $\mu=\tilde{\mu}\circ\pi^{-1}$. This $\mu$ is the law of unlabelled datapoints. Now restrict $\tilde{\mu}$ to the subspace $\Omega\times\{1\}$ and project this restriction under $\pi$ onto $\Omega$. Denote this measure $\mu_{1}$: $\mu_{1}(A)=\tilde{\mu}(A\times\{1\})$. Since $\mu_{1}$ is clearly absolutely continuous with regard to $\mu$, there exists for $\mu$-a.e. $x$ the Radon-Nykodym derivative:
$\eta(x)$ $=\frac{d\mu_{1}}{d\mu}(x)\in[0,1]$
$=P[Y=1\mid X=x].$
This is the conditional probability for a point $x$ to be labelled $1$. In this context, $\eta$ is referred to as the regression function. Since the measure $\tilde{\mu}$ is completely recoverable from the pair $(\mu,\eta)$, it is an often technically preferable way to encode the learning problem.
The learning (or generalization) error of a classifier, $T\colon\Omega\to\{0,1\}$ is the quantity
$\ell_{\tilde{\mu}}(T)$ $=\tilde{\mu}\{(x,y)\in\Omega\times\{0,1\}\colon T(x)\neq y\}$
$=P[T(X)\neq Y].$
The infimum of errors of all classifiers (which is in fact the minimum) is called the Bayes error:
$\ell^{\ast}_{\tilde{\mu}}$ $=\inf_{T\colon\Omega\to\{0,1\}}\ell_{\tilde{\mu}}(T)$
$=\int_{\Omega}\min\{\eta(x),1-\eta(x)\}\,d\mu(x).$
A classifier is called a Bayes classifier if its error equals the Bayes error. Those are exactly all classifiers satisfying $\mu$-a.e. the property
$T_{bayes}(x)=\begin{cases}1,&\mbox{ if }\eta(x)>\frac{1}{2},\\ 0,&\mbox{ if }\eta(x)<\frac{1}{2}.\end{cases}$
The Bayes error is zero if and only if the regression function is deterministic, that is, for $\mu$-a.e. $x$, $\eta(x)\in\{0,1\}$.
Given a learning rule $g$, the error of $g_{n}$ is a random variable:
$\ell_{\mu}(g_{n})=\ell_{\tilde{\mu}}(g_{n}(\sigma)),~~\sigma\sim\tilde{\mu}^{n}.$
The learning rule $g$ is called consistent under $\tilde{\mu}$ if the error of $g_{n}$ converges to the Bayes error in probability (equivalently, in expectation) as $n\to\infty$:
$\ell_{\tilde{\mu}}(g_{n})\overset{p}{\to}\ell^{\ast}_{\tilde{\mu}}$.
The rule $g$ is universally consistent on the measurable space $\Omega$ if it is consistent under every probability measure on the measurable space $\Omega\times\{0,1\}$.
Sometimes the consistent rules are called weakly consistent, to distinguish them from strongly consistent rules, where the convergence to the Bayes error is along a.e. every sample path. However, the strong consistency of learning rules appears to be a rare phenomenon (see [21], Subsection 2.3).
We are studying the $k$ nearest neighbour rule for supervised binary classification. The $k$-NN classifier assigns the label 0 or $1$ to every point of $\Omega$ according to the majority vote among the labels of the $k$ nearest neighbours in a random labelled $n$-sample. The nearest neighbours are chosen with regard to the metric $\rho$ on $\Omega$. There are two tie problems to resolve.
The first is the distance tie. Denote $r_{k}(x)=r_{k}(\sigma,x)$ the smallest radius of a closed ball around $x$ containing $k$ elements of the random sample $\sigma$:
$r_{k}(x)=\min\left\{r\geq 0\colon\sharp\{i\colon\rho(x,x_{i})\leq r_{k}\}\geq k\right\}.$
It may happen that there are strictly more than $k$ points at a distance $\leq r_{k}(x)$ from $x$. In this case, one needs an agreed algorithmic procedure to select the $k$ nearest neighbours of $x$: one selects all elements of the open $r_{k}$-sphere plus the necessary number of those sitting on the boundary. The usual choice is a random uniform tie-breaking. One chooses an auxiliary i.i.d. random sequence $\xi_{i}$ following some non-atomic probability distribution on the real line, independent of the data, and when having to choose between a few nearest neighbours at the distance $r_{k}(x)$ from $x$, one gives preference to the indices $i$ for which $\xi_{i}$ is smaller.
@@
The second kind of tie is the voting tie when $k$ is even and exactly half among the $k$ nearest neighbours of $x$ are labelled 0 and other half, $1$. Here the tie-breaking can be any, as it can be proved that it has no influence on the consistency of the $k$-NN classifier. Most often, one breaks the voting tie by choosing the label $1$.
The $k$-NN learning rule is a family of classifiers of the above form, under a choice of a sequence of values $k=k_{n}$ satisfying $k\to\infty$ and $k/n\to 0$ as $n\to\infty$.
@@
Now we will survey the previously known implications in Theorem 1.
@@ 2.3. Implication (2) $\Rightarrow$ (1)
$\lim_{r\downarrow 0}\frac{1}{\mu(B_{r}(x))}\int_{B_{r}(x)}f(x^{\prime})\,d\mu(x^{\prime})=f(x).$ (1)
@@
$\frac{1}{\mu(B_{r}(x))}\int_{B_{r}(x)}f(x^{\prime})\,d\mu(x^{\prime})\to f(x)\mbox{ as }r\downarrow 0,$ (2)
@@
$\mu\left\{x\in\Omega\colon\left|\frac{1}{\mu(B_{r}(x))}\int_{B_{r}(x)}f(x^{\prime})\,d\mu(x^{\prime})-f(x)\right|>{\epsilon}\right\}\to 0\mbox{ when }r\downarrow 0.$
The original theorem in [7] stated that for every complete separable metric space equipped with a probability measure $\mu$ and a regression function $\eta$ satisfying the weak Lebesgue–Besicovitch differentiation property the $k$-NN classifier is consistent.
However, one of the authors (Frédéric Cérou) has kindly pointed out to us [6] that the very last assertion in the proof of the main result of [7], at the bottom of p. 345, requires the strong Lebesgue–Besicovitch property. We reproduce the relevant part of the proof.
“Fix $\varepsilon>0$, then for every $\delta_{0}>0$
$P\left(\mu\left(B_{X,d_{k+1}}\right)^{-1}\int_{B_{X,d_{k+1}}}\left|\eta-\eta(X)\right|\,d\mu>\varepsilon\right)$ $\leq P\left(d_{(k+1)(X)}\geq\delta_{0}\right)+$
$\sup_{0\leq\delta\leq\delta_{0}}P\left(\left(\mu(B_{X,\delta}))\right)^{-1}\int_{B_{X,\delta}}\left|\eta-\eta(X)\right|\,d\mu>\varepsilon\right).$
Now, the first term goes to zero thanks to Cover and Hart’s result and the second one also thanks to the Besicovitch condition.”
Here $d_{k+1}$ is the radius of the smallest ball containing $k+1$ nearest neighbours of a random element $X$, in our notation it would be $r_{k+1}$. To better see the problem, let us rewrite the inequality measure-theoretically, denoting $d_{k+1}$ by $r(x)$:
$\mu\left\{x\colon\left|\frac{1}{\mu(B_{r(x)}(x))}\int_{B_{r(x)}(x)}f(x^{\prime})\,d\mu(x^{\prime})-f(x)\right|>\varepsilon\right\}$ $\leq\mu\left\{x\colon r\geq\delta_{0}\right\}+$
$\sup_{0\leq\delta\leq\delta_{0}}\mu\left\{x\colon\left|\frac{1}{\mu(B_{\delta}(x))}\int_{B_{\delta}(x)}f(x^{\prime})\,d\mu(x^{\prime})-f(x)\right|>\varepsilon\right\}.$
If $r(x)=d_{k+1}$ were a constant, it would be perfectly correct. But the radius depends on $x$, and so to guarantee the inequality, in the last term we must swap the supremum and the measure sign:
$\mu\left\{x\colon\left|\frac{1}{\mu(B_{r(x)}(x))}\int_{B_{r(x)}(x)}f(x^{\prime})\,d\mu(x^{\prime})-f(x)\right|>\varepsilon\right\}$ $\leq\mu\left\{x\colon r\geq\delta_{0}\right\}+$
$\mu\left\{x\colon\sup_{0\leq\delta\leq\delta_{0}}\left|\frac{1}{\mu(B_{\delta}(x))}\int_{B_{\delta}(x)}f(x^{\prime})\,d\mu(x^{\prime})-f(x)\right|>\varepsilon\right\}.$
In this form, the last term converges to zero when $\delta_{0}\to 0$ if (and only if) the function $\eta$ has the strong Lebesgue–Besicovitch property.
(This author must confess that this subtlety had escaped his attention during what he thought was a very careful reading of the paper [7].)
@@
[Cérou and Guyader [7]] Let $\Omega$ be a complete separable metric space equipped with a probability measure, $\mu$, and a $\mu$-measurable regression function $\eta\colon\Omega\to[0,1]$. If $\eta$ satisfies the strong Lebesgue–Besicovitch differentiation property with regard to $\mu$, then the $k$-NN classifier is consistent with the learning problem given by $(\mu,\eta)$ whenever $n,k\to\infty$ and $k/n\to 0$.